OpenStudy (spring98):

A closed container has 3.06 • 10^23 atoms of a gas. Each atom of the gas weighs 1.67 • 10^–24 grams. Which of the following shows and explains the approximate total mass, in grams, of all the atoms of the gas in the container?

OpenStudy (spring98):

0.47 grams, because (3.06 + 1.67) • (10^23 • 10^–24) = 4.73 • 10^–1 0.51 grams, because (3.06 • 1.67) • (10^23 • 10^–24) = 5.1102 • 10^–1 4.73 grams, because (3.06 + 1.67) • (10^23 • 10^–24) = 4.73 5.11 grams, because (3.06 • 1.67) • (10^23 • 10^–24) = 5.1102^

OpenStudy (spring98):

@satellite73

OpenStudy (spring98):

can you plz help me quick i only have 15 min. left.

OpenStudy (spring98):

@satellite73